An object is displaced from position vector r1 = (2+ 3j)m to r2 = (4+ 6j)m under a force = (3x2+ 2yj)N. Find the work down by this force.

Grade:12th pass

2 Answers

Sujit Kumar
5 years ago
NOTE : '[' is the integration symbol
 
W.D = 2[4 3xdx + 3[6 2y dy
= x2|+ y2 3|= 64 – 8 + 36 – 9 = 56 + 27 = 83J
Khimraj
5 years ago
Work done = \int_{x1}^{x2}F_{x}dx + \int_{y1}^{y2}F_{y}dy = \int_{2}^{4}3x^{2}dx + \int_{3}^{6}2ydy = 83J
Hope it clears.